OpenStudy (anonymous):

Rhombus ABCD is similar to the rhombus FGHR. If the difference between two distinct angles in ABCD is 48 degrees, then which is the measure of the larger angle in FGHR? Answer 114 degrees 112 degrees 102 degrees 94 degrees

OpenStudy (anonymous):

its 114 :)

OpenStudy (anonymous):

if the angle is A and the larger angle is A + 48 A + A + 48 = 180 A = (180 - 48)/2 = 66 and the larger angle is A + 48 = 66 + 48 = 114, the same in the other rhombus, which is similar